OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and that passes through the given point. y = –5x + 3; (–6, 3)

OpenStudy (anonymous):

y = 5x – 9 y = –5x + 9 y = –5x – 27 y = –5x + 27

OpenStudy (jjwalton04):

y = –5x – 27

OpenStudy (anonymous):

ty

OpenStudy (jjwalton04):

np